OpenStudy (anonymous):

The pH of a substance is 6. What is the molarity of the [H3O+]? Select one: a. 6.0 M b. 6.0 x 101 M c. 1.0 x X-6 M d. 1.0 x 106 M @Joel_the_boss @Kainui

JoelTheBoss (joel_the_boss):

Post this in chemistry not math...

OpenStudy (anonymous):

it has math in it

OpenStudy (gorv):

\[pH = - \log [H3O+].\]

OpenStudy (gorv):

pH=6

OpenStudy (anonymous):

So the answer is A

OpenStudy (gorv):

\[6= - \log [H3O+].\]

OpenStudy (gorv):

\[-6 = \log [H3O+].\]

OpenStudy (anonymous):

or is it C

OpenStudy (gorv):

log has base 10

OpenStudy (gorv):

now use this property

OpenStudy (gorv):

\[\log_ab=c\]

OpenStudy (gorv):

can be written as \[b=a^c\]

OpenStudy (gorv):

a= base and in Q log has base 10

OpenStudy (anonymous):

Isnt the asnwer C

OpenStudy (gorv):

yes it is